A corporation borrows $900,000 to expand its line of clothing. Some of the money is borrowed at 10%, some at 8% and at 9%. The annual interest payment to the lenders is $133,000. The amount borrowed at 8% is four times the amount borrowed at 10% . How much is borrowed at each rate ?

Can someone please help me solve this problem ?
Thank you so much.

Answer by stanbon(75887)

A corporation borrows $900,000 to expand its line of clothing. Some of the money is borrowed at 10%, some at 8% and at 9%. The annual interest payment to the lenders is $133,000. The amount borrowed at 8% is four times the amount borrowed at 10% . How much is borrowed at each rate ?
--------------
Equations:
Quantity Eq.::: x + y + z = 900,000 dollars
Interest Eq.:::0.10x + 0.08y +0.09z = 133000
Quantity Eq.:::y = 4x
---------------------------------
Substitute 4x for y:
x + 4x + z = 900,000
0.10x + 0.08(4x) + 0.09z = 133000
-----------------------------------------
Simplify:
5x + z = 900,000
0.42x+0.09z = 133000
----
Multiply thru the 1st Eq. by 42.
Multiply thru the 2nd Eq. by 500
-----
210x + 42z = 42*900,000
210x + 45z = 500*133000
----------------------------------
Comment: Those equation conditions result
in negative variable values.
Please check the wording of your problem.
Cheers,
Stan H.